### v1

# revtex4-1Repair the float

Abstract
This manuscript revisits the phenomenological emergent dark energy model (PEDE) by confronting it with recent cosmological data from early and late times. In particular we analyze PEDE model by using the baryon acoustic oscillation (BAO) measurements coming from both Dark Energy Spectroscopy Instrument (DESI) data release 1 and Sloan Digital Sky Survey (SDSS). Additionally, the measurements from cosmic chronometers, supernovae type Ia (Pantheon+), quasars, hydrogen II galaxies and cosmic background radiation distance priors are considered. By performing a Bayesian analysis based on Monte Carlo Markov Chain, we find consistent results on the constraints when SDSS and DESI are considered. However, we find higher values on the Hubble constant than Supernova $H_{0}$ for the Equation of State (SH0ES) does although it is still in agreement, within $1\sigma$ confidence level, when BAO measurements are added. Furthermore, we estimate the age of the Universe younger $\sim 3\%$ than the one predicted by the standard cosmology. Additionally, we report values of $q_{0}=-0.771^{+0.007}_{-0.007}$, $z_{T}=0.764^{+0.011}_{-0.011}$ for the deceleration parameter today and the deceleration-acceleration transition redshift, respectively.
pacs:
Dark energy, PEDE, cosmology
I Introduction
The description of our Universe is one of the most challenging topics in modern physics. This is because there are old unsolved problems, like the existence of dark matter (DM) related to the evolution of the structure in our Universe and the observed current acceleration Riess et al. (1998); Perlmutter et al. (1999), described by a de Sitter geometry and possibly caused by a dark energy (DE) fluid. In the DM case, we have many candidates, from supersymmetric particles to scalar fields Arbey and Mahmoudi (2021), while for DE we have modifications to Einstein Field equations to different fluids that mimics the accelerated evolution Motta et al. (2021); Herrera-Zamorano et al. (2020); Hernandez-Almada et al. (2019). However, the best candidates are a DM that behaves as cold fluid and a DE as the well known cosmological constant. Is in this vein, that the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) emerges, having unprecedented achievements in the understanding of the Universe evolution. Despite the success of the $\Lambda$CDM model, we do not have a full understanding of, for example, what the cosmological constant is. In this case, our best calculations predict an ultraviolet divergence when all the quantum vacuum fluctuations are calculated through the quantum field theory. This problem is known as the fine tuning problem (see Weinberg (1989); Zeldovich (1968)).
On the other hand, we have new challenges like those related to the $H_{0}$ and $\sigma_{8}$ tension detected by several observations Di Valentino et al. (2021a); Motta et al. (2021); Bamba et al. (2012); Di Valentino et al. (2021b, c, d, e); Perivolaropoulos and Skara (2022); Abdalla et al. (2022). Regarding the $H_{0}$ tension, this is due to the discrepancy between observations based on the Cosmic Microwave Background Radiations (CMB) by Planck Aghanim and et. al. (2020) and local observations given by Supernovae of the Ia type (SNIa) Riess et al. (2019). The disagreement is approximately at $\sim 5\sigma$ and, until now, we do not know if it is caused by systematic errors or it is due to mistaken theoretical assumptions. However, if the tension is confirmed, this event open a new window for possible new physics beyond the standard cosmological model. In addition, there is also tension in the amplitude of matter perturbations smoothed over $8h^{-1}$pc, known as $\sigma_{8}$, between the $\Lambda$CDM inferred value from CMB anisotropies and those obtained from weak gravitational lensing Abbott et al. (2022); Asgari et al. (2021); Heymans et al. (2021).
A natural and interesting alternative to resolve the problem of the Universe acceleration is based on the phenomenological emergent dark energy model (PEDE) proposed by Li and Shafieloo (2019). In this case, PEDE has the same parameters of $\Lambda$CDM, however, the causative of the Universe acceleration is a function that contains a tangent hyperbolic behavior with an inner logarithmic function. An interesting characteristic of the PEDE function is that it is negligible at early times, i.e. not affecting the physics of the early Universe, but at late times the contribution emerge producing the observed late time acceleration. Li and Shafieloo Li and Shafieloo (2019) constraint the model using SNIa, CMB and baryon acoustic oscillations (BAO), claiming a possible alleviation to the $H_{0}$ tension. Later on, in Pan et al. (2019) the authors show such alleviation within $68\%$ of the confidence level by using mainly CMB observational data. In the same vein, Koo et al. (2020) reconstruct the cosmic expansion from SNIa data using a non-parametric iterative smoothing method. Finally, in Hernández-Almada et al. (2020) the authors constrain PEDE using homogeneous and non-homogeneous Observational Hubble Data (OHD) samples. They report values consistent with $\Lambda$CDM and implement a robust dynamical system analysis, presenting the main stability points for this case.
Finally, we need to mention a recent paper Liu et al. (2024) where PEDE is studied with a barotropic fluid as a DM, called PEDE$+w_{dm}$. In this case, the authors implement constraints using CMB, BAO, SNIa with priors on $H_{0}$ and including recent datasamples Collaboration et al. (2024a, b); Bunker et al. (2024). Their results show, for the joint analysis, an increase of $H_{0}$ tension at $1.5\sigma$ but an alleviation of $S_{8}$ tension at $0.4\sigma$. Thus, they conclude that the PEDE$+w_{dm}$ is not a good alternative to the PEDE model.
The outline of the paper is as follows: Section II is dedicated to present a revision of the PEDE model and its description under a Friedmann-Lemaitre-Robertson-Walker cosmology. Section III contains the datasets used to constrain the model. In Section IV we present the results of our study, and finally in Section V we show our results and a discussion.
II Phenomenological emergent dark energy cosmology
In order to explore the PEDE cosmology, we use the Friedmann-Lemaitre-Robertson-Walker (FLRW) line element with flat curvature $ds^{2}=-dt^{2}+a(t)^{2}(dr^{2}+r^{2}d\Omega^{2})$, where $a(t)$ is the scale factor and $d\Omega^{2}$ is the solid angle. In this case, we assume a perfect fluid energy-momentum tensor $T_{\mu\nu}=pg_{\mu\nu}+(\rho+p)u_{\mu}u_{\nu}$, where $p$, $\rho$ and $u_{\mu}$ are the pressure, density and cuadri-velocities respectively, with only time dependence in order to maintain the cosmological principle.
Using the Einstein field equations, PEDE cosmology is described by the dimensionless Friedmann equation, which takes the form
$E(z)^{2}=\Omega_{0m}(z+1)^{3}+\Omega_{0r}(z+1)^{4}+\Omega_{DE}(z),$ (1)
where $E(z)\equiv H(z)/H_{0}$, $\Omega_{0m}$, $\Omega_{0r}$ are matter and radiation density parameters respectively, $z$ is the redshift and the mathematical structure of $\Omega_{DE}(z)$ is the DE density parameter associated with PEDE as proposed by Li and Shafieloo Li and Shafieloo (2019)
$\Omega_{DE}(z)=\Omega_{0DE}[1-\tanh(\log_{10}(z+1))].$ (2)
Notice that PEDE model has the same free parameters as $\Lambda$CDM cosmology. Thus, the deceleration parameter read as
$q(z)$ $=$ $\frac{z+1}{2E(z)^{2}}\Big{[}3\Omega_{0m}(z+1)^{2}+4\Omega_{0r}(z+1)^{3}$ (3)
$-\Omega_{0DE}\frac{{\rm sech}^{2}\Big{[}\frac{\ln(z+1)}{\ln(10)}\Big{]}}{\ln(10)(z+1)}\Big{]}-1.$
Additionally, it is possible to write the equation of state (EoS) only caused by the PEDE dynamics, which takes the form
$w(z)_{PEDE}=-\frac{1}{3\ln(10)}(1+\tanh[{\rm log}_{10}(z+1)])-1.$ (4)
We can write the effective EoS where all the densities contributes as
$w(z)_{eff}$ $=$ $\frac{1}{3}\Big{[}\frac{z+1}{E(z)^{2}}\Big{(}3\Omega_{0m}(z+1)^{2}+4\Omega_{0r}(z+1)^{3}$ (5)
$-\Omega_{0DE}\frac{{\rm sech}^{2}\Big{[}\frac{\ln(z+1)}{\ln(10)}\Big{]}}{\ln(10)(z+1)}\Big{)}-3\Big{]}\,,$
calculated through the formula $13w(z)_{eff}=2q(z)-$ 1 where $q(z)$ is the deceleration parameter.
III Datasets and constraints
The PEDE parameter phase-space is given by $\Theta=\{h,\Omega_{0b},\Omega_{0m}\}$, where $h$ is the dimensionless Hubble constant, and $\Omega_{0b}$ and $\Omega_{0m}$ are the baryon and matter densities, respectively. The aim is to constrain them using CMB distance priors Aghanim and et. al. (2020), cosmic chronometers (CC) Moresco et al. (2016); Jiao et al. (2023); Tomasetti et al. (2023), SNIa Scolnic et al. (2018); Brout et al. (2022), QSO Cao, Shuo et al. (2017), hydrogen II galaxies (HIIG) González-Morán et al. (2019, 2021) and BAO datasets by applying Monte Carlo Markov Chain (MCMC) tools through the Emcee package Foreman-Mackey et al. (2013) under Python environment. Hence MCMC runs are stopped up to achieve the convergence of the chains based on the autocorrelation function. Additionally, the priors considered are Uniform distributions over $h:[0.2,1]$ and $\Omega_{0m}:[0,1]$, and a Gaussian distribution centered at $\Omega_{0b}h^{2}=0.02202\pm 0.00046$ Aghanim and et. al. (2020). We define a baseline dataset as the combination of the CMB distance prior, CC, SNIa, QSO, HIIG datasets. Hence, the figure-of-merit for the baseline analysis is obtained by building a Gaussian log-likelihood given as
$-2\ln(\mathcal{L}_{\rm baseline})\varpropto\chi^{2}_{\rm baseline}=\sum_{i}\chi^{2}_{i}\,,$ (6)
where the sum runs over each $\chi^{2}$ according to the baseline dataset. Additionally, we add BAO data to this baseline. In the following, a brief description of each sample is presented.
III.1 Cosmic Microwave Background
The most recent CMB anisotropy data come from Planck 2018 legacy dataset release Aghanim and et. al. (2020) and provide us strong constraints on the cosmological model parameters. A way to extract such constraints without using a full perturbative analysis is available through the named distance prior, which gives effective information from the CMB power spectrum through the acoustic scale $l_{A}$, the shift parameter $R$ and $\Omega_{0b}h^{2}$. The parameter $l_{A}$ is related to the CMB temperature power spectrum in the transverse direction and generates variations of the peak spacing, $R$ characterizes the temperature power spectrum along the line-of-sight direction and modifies the heights of the peaks.
The figure-of-merit is built as
$\chi^{2}_{\rm CMB}=\vec{X}^{T}{\rm Cov}_{\rm CMB}^{-1}\vec{X}\,,$ (7)
where $\vec{X}$ is the difference vector between the theoretical and observed quantities
$\vec{X}$ $=\begin{pmatrix}R-1.7493\\ l_{A}-301.462\\ \Omega_{0b}-0.02239\end{pmatrix}$ (8)
and $\rm{Cov}_{\rm CMB}^{-1}$ is the inverse of the covariance matrix given by
${\rm Cov}_{\rm CMB}^{-1}$ $=\begin{pmatrix}93.2052939&-1.34477655&1634.16998\\ -1.34477655&0.160560491&5.05824338\\ 1634.16998&5.05824338&78905.7112\end{pmatrix}\times 10^{3}$ (9)
estimated using the CMB distance priors derived by Chen et al. (2019) based on $\omega$CDM model. The theoretical value of $R$ is estimated by
$R=\sqrt{\Omega_{0m}H_{0}^{2}}\,r(z^{*})\,,$ (10)
where $r(z^{*})$ is the comoving distance from the observer to the redshift at the photon decoupling epoch $z^{*}$ given by
$r(z^{*})=\frac{c}{H_{0}}\int_{0}^{z^{*}}\frac{dz^{\prime}}{E(z^{\prime})}\,,$ (11)
where $c$ is the speed of light.
III.2 Cosmic Chronometers
Cosmic chronometers is the name of a useful cosmological model independent dataset of 33 Hubble parameter measurements obtained from differential age tools applied to galaxies Moresco et al. (2016); Jiao et al. (2023); Tomasetti et al. (2023). These points are considered uncorrelated and cover a redshift $1.9650.07<z<$ 1.965, thus a $\chi^{2}$-function can be used in the form
$\chi^{2}_{{\rm CC}}=\sum_{i=1}^{33}\left(\frac{H_{th}(z_{i},\Theta)-H_{obs}(z_{i})}{\sigma^{i}_{obs}}\right)^{2},$ (12)
where the sum runs over each point in the dataset, $H_{th}(z_{i})$ is the theoretical Hubble parameter at the redshift $z_{i}$ and $(H_{obs})\pm\sigma_{obs}$ is the observational counterpart and its uncertainty, respectively.
III.3 Type Ia Supernovae
The Pantheon+ dataset Scolnic et al. (2018); Brout et al. (2022) is the largest sample of SNIa with a total of 1701 correlated distance modulus measurements which cover a redshift range of $2.260.001<z<$ 2.26. The theoretical distance modulus is defined by
$m_{th}=\mathcal{M}+5\log_{10}\left[\frac{d_{L}(z)}{10\,pc}\right],$ (13)
where $\mathcal{M}$ is a nuisance parameter and $d_{L}(z)$ is the luminosity distance estimated as
$d_{L}(z)=(1+z)\frac{c}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{E(z^{\prime})}.$ (14)
To avoid including this nuisance parameter in the statistics inference, it is convenient to use the following $\chi^{2}$ function Conley et al. (2010)
$\chi_{\rm SNIa}^{2}=a+\log\left(\frac{e}{2\pi}\right)-\frac{b^{2}}{e},$ (15)
where
$a$ $=$ $\Delta\boldsymbol{\tilde{\mu}}^{T}\cdot\mathbf{Cov_{P}^{-1}}\cdot\Delta\boldsymbol{\tilde{\mu}},$
$b$ $=$ $\Delta\boldsymbol{\tilde{\mu}}^{T}\cdot\mathbf{Cov_{P}^{-1}}\cdot\Delta\mathbf{1},$ (16)
$e$ $=$ $\Delta\mathbf{1}^{T}\cdot\mathbf{Cov_{P}^{-1}}\cdot\Delta\mathbf{1}\,,$
and $\Delta\boldsymbol{\tilde{\mu}}$ is the vector of the difference between the theoretical and observed distance modulus, $\Delta\mathbf{1}=(1,1,\dots,1)^{T}$ is the transpose of the unit vector and $\mathbf{Cov_{P}}$ is the covariance matrix.
III.4 Quasars
A sample of 120 distance modulus measurements of intermediate-luminosity quasars (QSO) spanning a redshift range $2.730.462<z<$ 2.73 is presented in Cao, Shuo et al. (2017). We constrain the cosmological parameters through the $\chi^{2}$ function
$\chi^{2}_{\rm QSO}=\sum_{i}^{120}\frac{[\theta_{th}(z_{i})-\theta_{obs}^{i}]^{2}}{\sigma^{2}_{i}},$ (17)
where $\theta_{i}\pm\sigma_{i}$ is the observed angular size and its $68\%$ confidence level uncertainty at the redshift $z_{i}$ and $\theta_{th}$ is the theoretical angular size, which expressed in terms of the angular diameter distance, $D_{A}(z)=d_{L}(z)/(1+z)^{2}$, is Sandage (1988)
$\theta(z)=\frac{l_{m}}{D_{A}(z)}\,,$ (18)
where $l_{m}$ is an intrinsic length fixed at the value $l_{m}=11.03\pm 0.25$ pc found by Cao, Shuo et al. (2017).
III.5 Hydrogen II Galaxies
A total of 181 distance modulus measurements coming from Hydrogen II galaxies spanning a redshift region $2.60.01<z<$ 2.6 are included in the Bayesian analysis González-Morán et al. (2019, 2021). These HIIG are compact galaxies with low mass with the characteristic that their luminosity is dominated by young massive burst of star formation Chávez et al. (2014). This allows to extract a correlation between the luminosity ($L$) and the inferred velocity dispersion ($\sigma$) of the ionized gas. In this case, the $\chi^{2}$-function can be written as
$\chi^{2}_{{\rm HIIG}}=\sum_{i}^{181}\frac{[\mu_{th}(z_{i},{\Theta})-\mu_{obs}(z_{i})]^{2}}{\epsilon_{i}^{2}}.$ (19)
The uncertainties of $\mu_{obs}$ at the redshift $z_{i}$ is $\epsilon_{i}$, being $\mu_{obs}$ the measured distance modulus obtained through
$\mu_{obs}=2.5(\alpha+\beta\log\sigma-\log f-40.08)\,,$ (20)
where $\alpha$ and $\beta$ are the intercept and slope of the $L$-$\sigma$ relation respectively, and $f$ is the measured flux.
The theoretical distance modulus is
$\mu_{th}(z,\Theta)=5\log_{10}\left[\frac{d_{L}(z)}{1\,{\rm Mpc}}\right]+25,$ (21)
where $d_{L}$ is the luminosity distance measured in Mpc.
III.6 Baryon Acoustic Oscillations
Baryon Acoustic Oscillation signal is a powerful tool to constrain cosmological parameters and is produced in the recombination epoch coming from the baryon-photon interactions. The typical measurements from BAO are the ratio $D_{V}(z)/r_{d}$ between the dilation scale $D_{V}(z)$ measured at the redshift $z$ and the size of the sound horizon at the drag epoch $r_{d}=r_{s}(z_{d})$, the ratio $D_{M}(z)/r_{d}$ where $D_{M}(z)$ is the comoving angular diameter, and $D_{H}(z)/r_{d}=(c/H(z))/r_{d}$. The dilation scale is defined as Eisenstein et al. (2005)
$D_{V}(z)=[z\,D_{H}(z)\,D_{M}^{2}(z)]^{1/3}\,,$ (22)
the sound horizon at $r_{d}$ is given by
$r_{d}=\int_{z_{d}}^{\infty}\frac{c_{s}(z)dz}{H(z)}\,,$ (23)
where $c_{s}(z)$ is the sound speed and
$D_{M}(z)=\frac{c}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{E(z^{\prime})}\,,$ (24)
is expressed for a flat cosmology. To confront the cosmological model to these measurements, we built a $\chi^{2}$ function of the form
$\chi^{2}_{\rm BAO}=\sum_{i}^{N}\left(\frac{O^{i}_{th}(z_{i})-O^{i}_{obs}(z_{i})}{\sigma_{i}}\right)^{2}$ (25)
where $O_{th}$ is the theoretical estimate of the observed counterpart $O_{obs}\pm\sigma_{i}$ at the redshift $z_{i}$, and the sum runs over the size of the sample. In this case, we will consider two samples, one comes from Sloan Digital Sky Survey (SDSS) and Dark Energy Survey (DES) and the second from the recent results from Dark Energy Spectroscopy Instrument (DESI).
III.6.1 BAO measurements from SDSS and DES
A total of 15 BAO measurements are used, 14 uncorrelated BAO quantities are reported in Alam et al. (2021) which cover a region $2.330.15<z<$ 2.33, and the recent BAO point $D_{M}(z_{eff})/r_{d}=19.51\pm 0.41$ at the effective redshift $z_{eff}=0.85$ reported by DES Collaboration et al. (2024a). We refer to this dataset as SDSS+DES.
III.6.2 BAO measurements from DESI
During the first year of DESI observations, the collaboration reported new 12 BAO measurements derived from BAO signatures in galaxy, quasars and Lyman-$\alpha$ forest tracers in the redshift range $4.160.1<z<$ 4.16 (see table 1 of Collaboration et al., 2024b). We use this sample to compare the parameter constraints with the SDSS dataset. We refer to this dataset as DESI-1Y.
IV Results
This section is devoted to present our results for the PEDE cosmology. Figure 1 shows the 2D contour distribution at 68% ($1\sigma$) confidence level (CL) and 99.7% ($3\sigma$) CL of the parameters, and the 1D posterior distribution using the baseline dataset (CMB+CC+SNIa+QSO+HIIG), baseline+SDSS+DES, and baseline+DESI-1Y datasets, respectively. Best fit parameters and their uncertainties at $1\sigma$ CL obtained using these datasets and only SDSS+DES and DESI-1Y data are shown in Table 1.
Table 1: Best fit values and their $1\sigma$ of uncertainties for PEDE cosmology using Baseline (CMB+CC+SNIa+QSO+HIIG), Baseline+SDSS+DES, Baseline+DESI-1Y. The last two row are the corresponding values when only SDSS+DES and DESI-1Y are considered respectively.
Parameter Baseline Baseline+SDSS+DES Baseline+DESI-1Y SDSS+DES DESI-1Y
PEDE
$h$ $0.761^{+0.005}_{-0.005}$ $0.739^{+0.004}_{-0.004}$ $0.743^{+0.004}_{-0.004}$
$\Omega_{0b}$ $0.040^{+0.001}_{-0.001}$ $0.042^{+0.0003}_{-0.0003}$ $0.041^{+0.0003}_{-0.0003}$ $0.034^{+0.012}_{-0.020}$ $0.034^{+0.012}_{-0.020}$
$\Omega_{0m}$ $0.242^{+0.005}_{-0.004}$ $0.264^{+0.004}_{-0.004}$ $0.259^{+0.004}_{-0.004}$ $0.299^{+0.019}_{-0.018}$ $0.296^{+0.016}_{-0.015}$
$\tau_{U}\,[\rm{Gyrs}]$ $13.514^{+0.018}_{-0.018}$ $13.570^{+0.017}_{-0.017}$ $13.560^{+0.017}_{-0.017}$
$z_{T}$ $0.811^{+0.013}_{-0.013}$ $0.753^{+0.011}_{-0.011}$ $0.764^{+0.011}_{-0.011}$ $0.665^{+0.043}_{-0.043}$ $0.673^{+0.036}_{-0.037}$
$q_{0}$ $-0.801^{+0.008}_{-0.008}$ $-0.764^{+0.007}_{-0.007}$ $-0.771^{+0.007}_{-0.007}$ $-0.703^{+0.033}_{-0.031}$ $-0.708^{+0.028}_{-0.026}$
$\chi^{2}$ $5860.78$ $5946.11$ $5922.22$ $21.09$ $16.28$
$\Lambda$CDM
$h$ $0.728^{+0.004}_{-0.004}$ $0.709^{+0.004}_{-0.004}$ $0.712^{+0.004}_{-0.004}$
$\Omega_{0b}$ $0.044^{+0.001}_{-0.001}$ $0.046^{+0.001}_{-0.001}$ $0.046^{+0.001}_{-0.001}$ $0.032^{+0.013}_{-0.019}$ $0.033^{+0.013}_{-0.019}$
$\Omega_{0m}$ $0.256^{+0.005}_{-0.005}$ $0.277^{+0.004}_{-0.004}$ $0.274^{+0.004}_{-0.004}$ $0.294^{+0.020}_{-0.018}$ $0.297^{+0.017}_{-0.016}$
$\tau_{U}\,[\rm{Gyrs}]$ $13.897^{+0.015}_{-0.015}$ $13.929^{+0.015}_{-0.015}$ $13.925^{+0.015}_{-0.015}$
$z_{T}$ $0.774^{+0.012}_{-0.012}$ $0.719^{+0.011}_{-0.011}$ $0.726^{+0.011}_{-0.011}$ $0.678^{+0.043}_{-0.044}$ $0.670^{+0.037}_{-0.038}$
$q_{0}$ $-0.778^{+0.008}_{-0.008}$ $-0.741^{+0.007}_{-0.007}$ $-0.746^{+0.007}_{-0.007}$ $-0.712^{+0.034}_{-0.031}$ $-0.706^{+0.029}_{-0.027}$
$\chi^{2}$ $5825.18$ $5908.01$ $5894.08$ $16.04$ $15.33$
$\Delta$AIC 35.59 38.09 28.14 5.05 0.95
Figure 1: 2D parameter distributions at $1\sigma$ (inner region) and $3\sigma$ (outermost region) for PEDE cosmology.
According to our results, the best fit value of $h$ is deviated $2.65\sigma$ from the one reported by SH0ES Riess et al. (2022) when the baseline is used, and this distance is reduced to $0.77\sigma$ ($1.13\sigma$) for baseline+SDSS+DES (baseline+DESI-1Y). Thus PEDE cosmology could alleviate the Hubble constant tension without adding more free parameters than those presented in $\Lambda$CDM, which is consistent with those results obtained by Li and Shafieloo (2019); Wang et al. (2024) using different samples. It is worth to mention that there is a degeneracy between $h$ and $r_{d}$ when only SDSS+DES data are used, consequently we do not give a value for $h$. With respect to matter density, we find higher values (around $13\%$ greater) for the single BAO samples than those using the rest of the data. This result is consistent with those reported in Li and Shafieloo (2019). However, for both SDSS+DES and DESI-1Y samples, we find that all quantities are in agreement for both PEDE and $\Lambda$CDM models (see Tab. 1 ). Regarding the age of the Universe $\tau_{U}$, we find that PEDE model predicts a younger Universe ($\sim 3\%$) when using baseline, baseline+SDSS+DES, and baseline+DESI-1Y than the ones predicted by $\Lambda$CDM. Notice that, due the degeneracy between $\tau_{U}$ and $h$, those values are not reported when only BAO data is used.
Figure 2 displays the Hubble parameter, deceleration parameter and the effective EoS in the redshift range $-0.95<z<2.2$ for baseline, baseline+SDSS+DES, and baseline+DESI-1Y datasets respectively. PEDE model predicts lower values of $H(z)$ in the region $z\gtrsim 0.2$ and higher values for the future than those estimated by the standard cosmology, indicating a higher expansion rate in the future. About the deceleration parameter, we extract its current value $q_{0}$ for each dataset and find that they deviate more than $3\sigma$ from those estimated from $\Lambda$CDM. Similarly, we obtain the corresponding values for the redshift of the deceleration-acceleration transition $z_{T}$, for which PEDE cosmology predicts an earlier transition than the standard cosmology. Additionally, $q(z)$ from both cosmologies converges in the future to the value $q(z\to-1)\to-1$. Regarding the $w_{eff}$, we observe that the PEDE cosmology behaves as a fluid that mimics a quintessence fluid in the past and as a phantom fluid in the future $z\lesssim-0.1$. Furthermore, in the region $z\gtrsim 0.7-0.8$, the effective EoS does not allow a Universe acceleration according to the restriction $w<-1/3$ imposed by general relativity, instead the Universe is going to the matter phase in which $w_{eff}\to 0$ as expected.
As a test, we compare our baseline results with those obtained by adding 14 new HIIG points reported in Llerena et al. (2023); de Graaff et al. (2023) and collected in Chávez et al. (2024) that covers a redshift region $7.432.41<z<$ 7.43. Five of them were measured using the James Webb Space Telescope (JWST) as part of the JWST Advanced Deep Extragalactic Survey Bunker et al. (2024). We do not find a significant deviation from our baseline results shown in Tab. 1.
Figure 2: Reconstruction of the Hubble (left panel), the deceleration parameters (middle panel), and the effective equation of state (right panel) for PEDE model in the redshift range $-0.95<z<2.2$ using baseline (top panel), baseline+SDSS+DES (middle panel) and baseline+DESI-1Y (bottom panel). Blue bands represent $3\sigma$ CL uncertainties of the PEDE cosmographic parameters. The dotted red line represents the evolution for the standard $\Lambda$CDM model. Baseline refers to the combined sample of CMB+CC+SNIa+QSO+HIIG.
[…]
V Summary and Discussions
PEDE cosmology is a successful model which describes the behavior of the DE component at late times by proposing an emergent DE, i.e. it is not an event that arise from the Big Bang as it happens with the cosmological constant, but instead the DE is described by a dynamical EoS and thus conserving the same degrees of freedom than the standard cosmology. PEDE cosmology has been confronted to several datasets previously, given interesting results regarding the Hubble tension, $S_{8}$ tension and others Abbott et al. (2022); Asgari et al. (2021); Heymans et al. (2021) (for excellent reviews of the $\Lambda$CDM status see Peebles (2024); Perivolaropoulos and Skara (2022)). For this reason, this work was devoted to revisit this emergent model by confronting it to recent data from SNIa (Pantheon+), QSO, HIIG, BAO (SDSS+DES and DESI), CC, and CMB distance prior. In this sense, the baseline data is defined as the sum of CMB, CC, SNIa, QSO, HIIG. Besides using the baseline, BAO and DESI-1Y databases by themselves, we included the combination of baseline+SDSS+DES and baseline+DESI-1Y. According to Table 1, the value of the Hubble constant is greater than the one from SH0ES Riess et al. (2022) when using baseline data but it moves towards the latter value when BAO measurements are added. Since the BAO measurements depend on the value of $r_{d}$ and there is a degeneracy between this value and $h$, both the values of $h$ and the age of the Universe, which is correlated with $h$, are not reported when only the BAO and DESI-1Y samples are considered individually. For matter and baryon densities, the results are in agreement with those obtained by the standard model, which is not the case for the redshift of the transition accelerated-decelerated Universe as PEDE cosmology predicts an earlier transition than the standard model. Similarly, the effective EoS has a trend to the $\Lambda$CDM EoS value at the future, and today ($z=0$) the $w_{eff}$ mimics a quintessence DE. Additionally, the age of the Universe $\tau_{U}$, we find that PEDE model predicts a younger Universe ($\sim 3\%$) when using baseline, baseline+SDSS+DES, and baseline+DESI-1Y than the ones predicted by $\Lambda$CDM.
Emergent DE models are the cornerstone to understand the current Universe acceleration without the need of a cosmological constant and its unsolvable problems. Mainly because the emergent models do not appear with the Big Bang as cosmological constant does, they manifest only at recent epochs and are produced by some unexplored events. In addition, PEDE in particular could be the key to alleviate the problem in $H_{0}$ as Ref. Pan et al. (2019) states, noticing a reduction within $60\%$ confidence level with an improvement of the $\chi^{2}$ for CMB. The key for this alleviation could be the presence of DE dynamics at $z\to 0$ instead the cosmological constant with its presence (almost in all parts negligible) since the beginning of the Universe. However, it is not yet clear which is the dynamical event producing the functional form of PEDE, being the best candidate a scalar field with the capability to reproduce Eq. (2) as Ref. Mortonson et al. (2009) argue.
Acknowledgements.
A.H.A. thanks to the support from Luis Aguilar, Alejandro de León, Carlos Flores, and Jair García of the Laboratorio Nacional de Visualización Científica Avanzada. M.A.G.-A. acknowledges support from cátedra Marcos Moshinsky (MM), Universidad Iberoamericana for support with the SNI grant and the numerical analysis was also carried out by Numerical Integration for Cosmological Theory and Experiments in High-energy Astrophysics (Nicte Ha) cluster at IBERO University, acquired through cátedra MM support. A.H.A, V.M. and M.A.G.-A acknowledge partial support from project ANID Vinculación Internacional FOVI220144. M.L.M.M. thanks to the SNII-Mex́ico, CONAHCyT for support.